The length of the hypotenuse of a 30-60-90 triangle is 13 meters. Find the length of the opposite side 30 degree angle.
Answer by CharlesG2(828)

The length of the hypotenuse of a 30-60-90 triangle is 13 meters. Find the length of the opposite side 30 degree angle.
sohcahtoa: sin = opposite(opp)/hypotenuse(hyp),
cos = adjacent(adj)/hyp,
tan = opp/adj
a 30-60-90 triangle is half an equilateral triangle
if the equilateral triangle has sides 1,1, and 1 then:
sin 30 = 0.5/1 = 0.5 (0.5 is half of 1)
cos 30 = adj/1 --> adj = sqrt(3)/2
30-60-90 triangles have the ratio 0.5:sqrt(3)/2:1 as their sides
so if the hypotenuse is 13 meters then the side opposite the 30 degree angle is:
13 * 0.5 = 6.5 meters